a clerk cut a rope into 8 pieces of equal length. If each piece is 3 and1/2 feet long, how long was the rope before it was cut?
Answer by oberobic(2304)

8 pieces @ 3 1/2 ft per piece = 8* 7/2 = 56/2 = 28 ft.